Hello,
I am stuck again and at the end of my rope, please help me!! I need help with the following problems:
1) Find the equation, in standard form, of the line perpendicular to 2x + 3y = -5 and passing through (3,-5). Write the equation in standard form, with all integer coefficients.
2) Solve the system of equations using the substitution method. If the answer is a unique solution, present it as an ordered pair: (x,y). If not, specify whether the answer is "no solution" or "infinitely many solutions" and state how you arrived at that conclusion.
-7x + y = 4
2x + 8y = 0
Any help you can give will be greatly appreciated.
Thanks

Answer by Mathtut(3670)

1) Find the equation, in standard form, of the line perpendicular to 2x + 3y = -5 and passing through (3,-5). Write the equation in standard form, with all integer coefficients
:
slope of 2x+3y=5 is 3y=-2x+5--->y=(-2/3)x+5........-2/3
:
slope of line perpendicular to this line would be -3/2
:
now using point slope formula y+5=-3/2(x-3)
:
y+5=-3/2x+9
:
2y+10=-3x+28
:
3x+2y=18(standard form)